OpenStudy (anonymous):

Yesterday, Maria went on a bike ride. Her average speed was 10 miles per hour. Today, she went on another ride, this time averaging 14 miles per hour. In the two days, she biked for a combined total time of 10 hours. Let x be the number of hours she biked yesterday. Write an expression for the combined total number of miles she biked in the two days.

OpenStudy (alexandervonhumboldt2):

s=v*t where s is distance, v is speed and t is time. 10*x+14*(10-x)=S s=140-4x